# Unveiling the optimization process of Physics Informed Neural Networks: How accurate and competitive can PINNs be?

### v2

@@ Abstract
We suggest adjustments to the BFGS optimization algorithm and the mean squared error loss function, which could significantly improve precision by several orders of magnitude.
@@
We demonstrate that our scheme is applicable to a wide range of physical problems, yielding more accurate results while reducing the computational cost.
@@ 1 Introduction
Recent advancements in Physics-Informed Neural Networks (PINNs) have positioned them as serious contenders in the domain of computational physics [1, 2], disrupting the longstanding monopoly held by classical numerical methods in many and varied physical applications, such as fluid and solid mechanics [3, 4], quantum optics [5], black-hole spectroscopy [6], or radiative transfer [7], among many others. This disruptive potential arises from their innate ability to integrate domain-specific physics principles with the powerful learning capabilities of neural networks.
@@
One of the most challenging aspects of neural networks is the inherently non-convex nature of their optimization problems. As a result, an increasing number of studies are delving deeper into their learning dynamics (see for example [8, 9, 10, 11]). Focusing specifically on PINNs, [12] analyses the convergence process using gradient-descent methods by examining the eigenvalue spectra of the Neural Tangent Kernel (NTK). The NTK has been discussed in many subsequent studies to provide a better understanding of the dynamics of the training process in many physical problems under different modifications of the standard PINN algorithm [13, 14, 15, 16]. The learning dynamics of PINNs has recently been studied in [17] through the lens of the gradient signal-to-noise ratio (GSNR), which let the authors characterize the optimization process of PINNs in different phases, as predicted by the Information Bottleneck theory [18]. As a complementary path, other works in the PINN literature have focused more on the development of new optimization techniques. In [19] they analyzed the behavior of gradient-descent based optimization methods, concluding that training with these algorithms is limited because of the ill-conditioning nature of the parameter space, and suggested various strategies to precondition this space. [20] introduced a novel second-order optimizer that significantly improves the solution returned by L-BFGS for some problems. Learnable optimization, also known in the field of deep learning as learning to learn [21] has also been applied recently to PINNs in [22], improving the results obtained with standard machine learning optimizers.
@@
While PINNs exhibit remarkable adaptability to complex physical systems and have demonstrated interesting results across various domains, their efficacy relies on several factors that warrant closer scrutiny. One critical aspect is the architectural design of the neural network [23, 24, 25, 26], which influences its ability to capture intricate physical phenomena accurately. Other ideas proposed to enhance the performance of PINNs are for example domain decompositions [27], trainable activation functions [28], loss function redefinitions [29], or curriculum strategies [30], among others.
@@
Despite their flexibility, training neural networks can be computationally expensive for many problems. The challenge also makes it difficult to determine which hyperparameters are crucial for improving results. Recent research has focused on this issue, particularly on the weights controlling the residual and supervised terms, with new methods proposed to mitigate the negative effects of their magnitude disparity [31, 32, 33]. Some studies have also tried to give estimates of the errors as a function of different hyperparameters, such as the network size or the number of training points, for particular PDE problems [34, 35, 36, 37, 38, 39]. Scaling PINNs to large-scale or high-dimensional problems can be challenging. The neural network may require more layers or neurons, increasing the risk of overfitting or making the training process unstable. Tackling the so-called Curse of Dimensionality is not a trivial task, and it has been recently studied in [40].
@@
The paper is structured as follows: Section 2 provides a brief overview of our PINN framework and discusses key issues related to commonly used optimizers. We dive into the details behind the optimization process in PINNs and we present our proposed contributions to improve it in Section 3. In Section 4, we thoroughly examine a relatively simple case to demonstrate the significant impact of selecting the appropriate optimizer and how such a choice can minimize network size while achieving excellent results, surpassing previous studies with similar problems. Additionally, in Section 5, we present a comprehensive set of physical problems spanning various fields, illustrating that the insights from the preceding section can be extended to a large variety of problems. Finally, Section 6 summarizes our findings and outlines potential paths for future improvements.
@@
where $k=2\pi/L$, one can indeed check that $u$ has the desired behaviour at all the borders. This parametrization enforces periodicity for $u$ and all its derivatives with respect to $x_{\gamma}$, provided that $f_{b}$ and all its derivatives are also periodic. However, this does not necessarily work for all cases (for example, if we want a periodic function but not its derivatives). For some problems, the full Fourier expansion is formally needed to compute the solution (see e.g. [57, 24, 58]). Another possible approach is to consider a different set of functions, such as Hermite-based interpolation polynomials [56], chosen to be periodic up to the desired order. These polynomials contain additional trainable parameters to enforce the appropriate relation between the function and the derivatives at the boundaries.
@@ 4.1 Current-free Grad-Shafranov equation (CFGS)
which is completely determined by the surface boundary conditions (providing the $b_{l}$ coefficients). This is a relatively simple problem to solve but, nevertheless, valuable conclusions can be drawn for the optimization procedure by analyzing it and comparing it with the analytical solution. The complexity of the problem depends on the number of multipoles considered in (32), i.e on the number of non-zero coefficients $b_{l}$. Simpler solutions will achieve the same accuracy with fewer trainable parameters compared to more complex ones. To illustrate this, we focus on a dipole-quadrupole solution ($b_{1},b_{2}\neq 0,b_{l>2}=0$). Subsequently, we will extend our findings to encompass a broader range of force-free solutions and various problem types.
@@ 4.1.1 Impact of the optimization algorithm
In a similar light, we can examine the impact of utilizing the modified versions of the loss function function discussed in section 3.3. To this end, we train three identical networks with different loss functions, namely $\mathcal{J},\mathcal{J}_{1/2},\mathcal{J}_{\log}$. We use the standard BFGS algorithm in order to isolate the impact of the loss function modifications from that of the BFGS modifications.
@@ 5 Other physics problems
Table 4: Architecture and training hyperparameters for various physical applications consdidered in section 5. In all cases, we use a $\tanh$ activation function for the hidden layers. The Neurons column refers to neurons per hidden layer. The Adam it. column refers to the number of iterations where the Adam optimizer is used before switching to a quasi-Newton method. The Batch size column refers to the number of points sampling the domain for each training set. The training set changes every 500 iterations in order to sample as many points as possible. The Iterations column denote the number of Quasi-Newton iterations.
@@
where $\mathcal{N}$ is the output of the network. The loss function is then constructed using equation (2), where $\mathcal{L}=\nabla^{2}$ and $G=-k^{2}u+q(x,y)$. We consider a low wavenumber case with $a_{1}=1$, $a_{2}=4$ and $k=1$, as in [57]. We remark that, if we consider periodic boundary conditions, the constant $k$ should be chosen such that the solution of the homogeneous PDE with periodic boundary conditions is zero. It can be readily seen that the homogeneus solution is zero, provided that $k^{2}\neq\pi^{2}\left(n^{2}+m^{2}\right)$, where $n,m\in\mathbb{Z}$.
@@
Then, the functions $f_{b}$ and $h_{b}$ used in (3) can be defined through the the following ansatz
@@
Results for the loss function are shown in figure 22. Figure 23 shows the solution for the components and the modulus of the velocity aquired with the self-scaled Broyden algorithm. Since no analytical or high-precision numerical solution is available, we estimate the error according to equation (91) as $\epsilon_{\mathrm{NN}}\approx 1.3\times 10^{-4}$, obtained with the self-scaled Broyden algorithm. The standard BFGS algorithm on the other hand achieves, if trained against the standard MSE loss $\mathcal{J},$ a value of $\epsilon_{\mathrm{NN}}\approx 1.0\times 10^{-3}$, whereas if the logarithm of the MSE loss $\mathcal{J}_{\mathrm{log}}$ is employed, this value is reduced up to $\epsilon_{\mathrm{NN}}\approx 1.6\times 10^{-4}$.
@@ 5.1 Comparison with the literature
In the current landscape of PINN research, L-BFGS has become a widely used optimizer due to its generally superior performance compared to the Adam optimizer. However, it has demonstrated some limitations, especially when confronted with highly ill-conditioned problems, as evidenced in some studies [105]. Consequently, it needs a significantly higher number of trainable parameters than the BFGS algorithm to achieve comparable precision. Although training two identical networks with BFGS would be markedly slower compared to L-BFGS, primarily due to the former’s requirement to store and update the inverse Hessian estimate at each iteration, with BFGS one can use much smaller networks with higher accuracy.
@@ Appendix B Efficient computation of the scaling parameter $\tau_{k}^{(1)}$
Note that the original suggestion of [74] for the scaling parameter was $\tau_{k}^{(1)}=\frac{-\bm{y}_{k}\cdot\bm{s}_{k}}{\alpha_{k}\bm{s}_{k}\cdot\nabla\mathcal{J}\left(\bm{\Theta}_{k}\right)}$, which was indeed motivated by the need to reduce the condition number of (35). However, this choice has shown to be inferior in terms of performance compared to the standard BFGS, when combined with an inexact line search computation of $\alpha_{k}$, as shown in [108] and was confirmed by our own analysis. Instead, we followed the suggestion introduced in [69], which is a simple modification of the original one, but ensures theoretically super-linear convergence with inexact line searches. It can be thought of as a switch between the standard BFGS algorithm ($\tau_{k}=1$) and the self-scaled BFGS algorithm of [74]. Curiously, we found that in the majority of the training iterations, the value of $\tau_{k}=1$ is selected, with only a small portion of them benefitting from the scaling. Still, this was sufficient to produce considerable improvements in the optimization process.

### v3

@@ Abstract
We suggest adjustments to the BFGS algorithm and MSE loss to greatly enhance precision by orders of magnitude.
@@
We show that our scheme applies to various physical problems, offering greater accuracy and lower computational cost.
@@ 1 Introduction
Recent advances in physics-informed neural networks (PINNs) have positioned them as serious contenders in the domain of computational physics [1, 2], disrupting the longstanding monopoly held by classical numerical methods in many and varied physical applications, such as fluid and solid mechanics [3, 4], quantum optics [5], black-hole spectroscopy [6], or radiative transfer [7], among many others. This disruptive potential arises from their innate ability to integrate domain-specific physics principles with the powerful learning capabilities of neural networks.
@@
One of the most challenging aspects of neural networks is the inherently non-convex nature of their optimization problems. As a result, an increasing number of studies are delving deeper into their learning dynamics (see, for example, [8, 9, 10, 11]). Focusing specifically on PINNs, [12] analyzes the convergence process using gradient descent methods by examining the eigenvalue spectra of the Neural Tangent Kernel (NTK). The NTK has been discussed in many subsequent studies to provide a better understanding of the dynamics of the training process in many physical problems under different modifications of the standard PINN algorithm [13, 14, 15, 16]. The learning dynamics of PINNs has recently been studied in [17] through the lens of the gradient signal-to-noise ratio (GSNR), which allowed the authors to characterize the optimization process of PINNs in different phases, as predicted by the Information Bottleneck theory [18]. As a complementary path, other works in the PINN literature have focused more on the development of new optimization techniques. In [19] they analyzed the behavior of gradient descent-based optimization methods, concluding that training with these algorithms is limited because of the ill-conditioning nature of the parameter space, and suggested various strategies to precondition this space. [20] introduced a novel second-order optimizer that significantly improves the solution returned by L-BFGS for some problems. Learnable optimization, also known in the field of deep learning as learning to learn [21] has also been applied recently to PINNs in [22], improving the results obtained with standard machine learning optimizers.
@@
Although PINNs exhibit remarkable adaptability to complex physical systems and have demonstrated interesting results across various domains, their efficacy is based on several factors that warrant closer scrutiny. One critical aspect is the architectural design of the neural network [23, 24, 25, 26], which influences its ability to capture intricate physical phenomena accurately. Other ideas proposed to enhance the performance of PINNs are, for example, domain decompositions [27], trainable activation functions [28], loss function redefinitions [29], or curriculum strategies [30], among others.
@@
Despite their flexibility, training neural networks can be computationally expensive for many problems. The challenge also makes it difficult to determine which hyperparameters are crucial for improving results. Recent research has focused on this issue, particularly the weights controlling the residual and supervised terms, with new methods proposed to mitigate the negative effects of their magnitude disparity [31, 32, 33]. Some studies have also tried to give estimates of the errors as a function of different hyperparameters, such as the network size or the number of training points, for particular PDE problems [34, 35, 36, 37, 38, 39]. Scaling PINNs to large-scale or high-dimensional problems can be challenging. The neural network may require more layers or neurons, increasing the risk of overfitting or making the training process unstable. Tackling the so-called Curse of Dimensionality is not a trivial task, and it has been recently studied in [40].
@@
The paper is structured as follows: Section 2 provides a brief overview of our PINN framework and discusses key issues related to commonly used optimizers. We dive into the details behind the optimization process in PINNs and we present our proposed contributions to improve it in Section 3. In Section 4, we thoroughly examine a relatively simple case to demonstrate the significant impact of selecting the appropriate optimizer and how such a choice can minimize network size while achieving excellent results, surpassing previous studies with similar problems. Additionally, in Section 5, we present a comprehensive set of physical problems spanning various fields, illustrating that the insights from the preceding section can be extended to a large variety of problems. Finally, Section 6 summarizes our findings and outlines potential paths for future improvements. All code accompanying this manuscript is available on GitHub at https://github.com/jorgeurban/self_scaled_algorithms_pinns.
@@
where $k=2\pi/L$, one can indeed check that $u$ has the desired behavior at all the borders. This parametrization enforces periodicity for $u$ and all its derivatives with respect to $x_{\gamma}$, provided that $f_{b}$ and all its derivatives are also periodic. However, this does not necessarily work for all cases (for example, if we want a periodic function but not its derivatives). For some problems, the full Fourier expansion is formally needed to compute the solution (see e.g. [57, 24, 58]). Another possible approach is to consider a different set of functions, such as Hermite-based interpolation polynomials [56], chosen to be periodic up to the desired order. These polynomials contain additional trainable parameters to enforce the appropriate relation between the function and the derivatives at the boundaries.
@@ 4.1 Current-free Grad-Shafranov equation (CFGS)
which is completely determined by the surface boundary conditions (providing the $b_{l}$ coefficients). This is a relatively simple problem to solve but, nevertheless, valuable conclusions can be drawn for the optimization procedure by analyzing it and comparing it with the analytical solution. The complexity of the problem depends on the number of multipoles considered in (32), i.e. on the number of non-zero coefficients $b_{l}$. Simpler solutions will achieve the same accuracy with fewer trainable parameters compared to more complex ones. To illustrate this, we focus on a dipole-quadrupole solution ($b_{1},b_{2}\neq 0,b_{l>2}=0$). Subsequently, we will extend our findings to encompass a broader range of force-free solutions and various problem types.
@@ 4.1.1 Impact of the optimization algorithm
In a similar light, we can examine the impact of utilizing the modified versions of the loss function discussed in section 3.3. To this end, we train three identical networks with different loss functions, namely $\mathcal{J},\mathcal{J}_{1/2},\mathcal{J}_{\log}$. We use the standard BFGS algorithm in order to isolate the impact of the loss function modifications from that of the BFGS modifications.
@@ 5 Other physics problems
Table 4: Architecture and training hyperparameters for various physical applications considered in section 5. In all cases, we use a $\tanh$ activation function for the hidden layers. The Neurons column refers to neurons per hidden layer. The Adam it. column refers to the number of iterations where the Adam optimizer is used before switching to a quasi-Newton method. The Batch size column refers to the number of points sampling the domain for each training set. The training set changes every 500 iterations in order to sample as many points as possible. The Iterations column denote the number of Quasi-Newton iterations.
@@
where $\mathcal{N}$ is the output of the network. The loss function is then constructed using equation (2), where $\mathcal{L}=\nabla^{2}$ and $G=-k^{2}u+q(x,y)$. We consider a low wavenumber case with $a_{1}=1$, $a_{2}=4$ and $k=1$, as in [57]. We remark that, if we consider periodic boundary conditions, the constant $k$ should be chosen such that the solution of the homogeneous PDE with periodic boundary conditions is zero. It can be readily seen that the homogeneous solution is zero, provided that $k^{2}\neq\pi^{2}\left(n^{2}+m^{2}\right)$, where $n,m\in\mathbb{Z}$.
@@
Then, the functions $f_{b}$ and $h_{b}$ used in (3) can be defined through the following ansatz
@@
Results for the loss function are shown in figure 22. Figure 23 shows the solution for the components and the modulus of the velocity acquired with the self-scaled Broyden algorithm. Since no analytical or high-precision numerical solution is available, we estimate the error according to equation (91) as $\epsilon_{\mathrm{NN}}\approx 1.3\times 10^{-4}$, obtained with the self-scaled Broyden algorithm. The standard BFGS algorithm on the other hand achieves, if trained against the standard MSE loss $\mathcal{J},$ a value of $\epsilon_{\mathrm{NN}}\approx 1.0\times 10^{-3}$, whereas if the logarithm of the MSE loss $\mathcal{J}_{\mathrm{log}}$ is employed, this value is reduced up to $\epsilon_{\mathrm{NN}}\approx 1.6\times 10^{-4}$.
@@ 5.1 Comparison with the literature
In the current landscape of PINN research, L-BFGS has become a widely used optimizer due to its generally superior performance compared to the Adam optimizer. However, it has demonstrated some limitations, especially when confronted with highly ill-conditioned problems. Consequently, it needs a significantly higher number of trainable parameters than the BFGS algorithm to achieve comparable precision. Although training two identical networks with BFGS would be markedly slower compared to L-BFGS, primarily due to the former’s requirement to store and update the inverse Hessian estimate at each iteration, with BFGS one can use much smaller networks with higher accuracy.
@@ Appendix B Efficient computation of the scaling parameter $\tau_{k}^{(1)}$
Note that the original suggestion of [74] for the scaling parameter was $\tau_{k}^{(1)}=\frac{-\bm{y}_{k}\cdot\bm{s}_{k}}{\alpha_{k}\bm{s}_{k}\cdot\nabla\mathcal{J}\left(\bm{\Theta}_{k}\right)}$, which was indeed motivated by the need to reduce the condition number of (35). However, this choice has shown to be inferior in terms of performance compared to the standard BFGS, when combined with an inexact line search computation of $\alpha_{k}$, as shown in [107] and was confirmed by our own analysis. Instead, we followed the suggestion introduced in [69], which is a simple modification of the original one, but ensures theoretically super-linear convergence with inexact line searches. It can be thought of as a switch between the standard BFGS algorithm ($\tau_{k}=1$) and the self-scaled BFGS algorithm of [74]. Curiously, we found that in the majority of the training iterations, the value of $\tau_{k}=1$ is selected, with only a small portion of them benefiting from the scaling. Still, this was sufficient to produce considerable improvements in the optimization process.